# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v3

@@ Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals in datasets with complete or missing data. SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation. Extensive empirical evaluations across 11 tabular datasets show that SEMF often achieves narrower normalized prediction intervals and higher coverage rates than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper. Furthermore, SEMF can be integrated with machine learning models like gradient-boosted trees and neural networks, highlighting its practical applicability. The results indicate that SEMF enhances uncertainty quantification, particularly in scenarios with complete data.
@@ 1 Introduction
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic). SEMF generates representations for latent or missing modalities, which can be relevant for incomplete data and holds potential for multi-modal data applications, though multi-modal settings are left for future exploration. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. SEMF differs from prominent supervised EM approaches such as [8], which focus on point prediction using Gaussian Mixture Models (GMMs). Additionally, our method operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions. Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper.
The remainder of this paper is organized as follows: Section 2 details the theory and the methodology of SEMF. Section 3 reviews related works in latent representation learning, uncertainty estimation, and handling of missing data. Section 4 describes the experimental setup, including datasets and evaluation metrics. Section 5 discusses the results, demonstrating the efficacy of SEMF. Lastly, Section 6 concludes the paper, and Section 7 outlines the limitations and potential research directions.
@@ 2 Method
This section presents the founding principles of SEMF from its parameters, training, and inference procedure with, at its core, the EM algorithm. This algorithm, first introduced by [5], is an unsupervised method for handling latent variables and incomplete data. Invented to maximize the model likelihood, it builds a sequence of parameters that guarantee an increase in the log-likelihood [9] by iterating between the Expectation (E) and the Maximization (M) steps. In the E-step, one computes
@@
where $p^{\prime}$ stands for the current estimates, $\log p(x,z)$ is the log-likelihood of the complete observation $(x,z)$, and $z$ is a latent variable. The M-step maximizes this $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
@@ 2.1 Problem Scenario
Let $x=(x_{1},x_{2},\ldots,x_{K})$ denote $K$ inputs and the output be $y$. For simplicity, we limit $y$ to be numerical, although it could be categorical without loss of generality. Component $x_{k}$ is a source: a modality, a single or group of variables, or an unstructured input such as an image or text. For clarity, we limit to $K=2$, where $x_{1}$ and $x_{2}$ are single variables. We assume that only $x_{1}$ may contain missing values, either at random or partially at random.
@@
where $p^{\prime}$ is our current estimate. Eq. 2 can be estimated by MC sampling. Since sampling from $p^{\prime}(z|y,x)$ can be inefficient, we rather rely on the decomposition $p^{\prime}(z|y,x)=p^{\prime}(y|z)p^{\prime}(z|x)/p^{\prime}(y|x)$. Thus, we sample $z_{r}$ from $p^{\prime}(z|x)$, $r=1,\ldots,R$, and, setting $w_{r}=p^{\prime}(y|z_{r})/\sum_{t}p^{\prime}(y|z_{t})$, approximate the right-hand side term of Eq. 2
@@
If $x_{1}$ is missing, a similar development leads to
[…]
where $x_{1,r}$ and $z_{r}$ are respectively sampled from $p^{\prime}(x_{1}|x_{2})$ and $p^{\prime}(z|x_{1,r},x_{2})$.
@@ 2.2 Objective Function
Adapting Eq. 3 and Eq. 4 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
$\mathcal{L}(\phi,\theta,\xi)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i,r})+\log p_{\theta}(y_{i}|z_{i,r})+\mathds{1}_{\{i\in I_{m}\}}\log p_{\xi}(x_{1,i,r}|x_{2,i})\right\}w_{i,r},$ (5)
where $I_{m}$ is the set of those $i$’s such that $x_{1,i}$ is missing. The models of $p(y|z)$, $p(z|x)$, and $p(x_{1}|x_{2})$ inherit parameters $\theta$, $\phi$, and $\xi$, respectively. Also, $x_{1,i,r}$ is sampled from $p_{\xi^{\prime}}(x_{1}|x_{2,i})$, if $x_{1,i}$ is missing, and $z_{1,i,r}$ and $z_{2,i,r}$ are sampled from $p_{\phi^{\prime}}(z_{1}|x_{1,i,r})$ and $p_{\phi^{\prime}}(z_{2}|x_{2,i})$. Furthermore, for compactness of notation, $x_{i,r}$ is $(x_{1,i},x_{2,i})$ if $x_{1,i}$ is observed, and $(x_{1,i,r},x_{2,i})$ if $x_{1,i,r}$ is missing. Finally, the weights are
$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}.$ (6)
Eq. 5 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, the decoder model, $p_{\theta}$, from the latent variables to the output, and, if applicable, the model handling missing data, $p_{\xi}$. At each M-step, ${\cal L}$ is minimized with respect to $\theta$, $\phi$, and $\xi$. Then, $\theta^{\prime}$, $\phi^{\prime}$, and $\xi^{\prime}$ are updated, as well as the weights and the sampling. Then the process is iterated until convergence.
2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
Similar to [7], we develop further ${\cal L}$ under normality assumptions for the encoder, $p_{\phi}$, and the decoder, $p_{\theta}$. These simple cases are illustrative, though any other distributions could be adopted, including non-continuous or non-numerical outcomes.
@@ 2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{1}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (8)
The mean $g_{\phi_{k}}(x_{k})$ can be any model, such as a neural network, with output of length $m_{k}$, $k=1,2$. The scale $\sigma_{k}$ can be fixed, computed via the weighted residuals, or learned through a separate set of models. It controls the amount of noise introduced in the latent dimension and is pivotal in determining the prediction interval width for $p(y|z)$. In this paper, $\sigma_{k}$ is fixed for simplicity.
@@
Model for missing data $p_{\xi}(x_{1}|x_{2})$.
We use an empirical model for $X_{1}$ given $X_{2}=x_{2}$ where $p_{\xi}(x_{1}|x_{2})$ put masses only on those non-missing $x_{1}$’s in the training set. Let $I_{nm}$ be the set of indices such that $x_{1,j}$, $j\in I_{nm}$, are all the non-missing $x_{1}$ in the training set. Additionally, for a given $j\in I_{nm}$, $x_{1}[j]$ is the observed $x_{1}$ corresponding to $j$. For a given $x_{2}$, $p_{\xi}(x_{2})$ is a vector of length $|I_{nm}|$, the cardinality of $I_{nm}$, with components
$p_{\xi}(j|x_{2})=\frac{\exp\{h_{\xi,j}(x_{2})\}}{\sum_{t\in I_{nm}}\exp\{h_{\xi,t}(x_{2})\}},\quad j\in I_{nm},$ (11)
where $h_{\xi}$ is a vector of length $|I_{nm}|$, typically a neural network with input $x_{2}$ and output on $I_{nm}$. Now, the probability of $X_{1}=x_{1}$ given $X_{2}=x_{2}$ is
$p_{\xi}(x_{1}|x_{2})=\sum_{j\in I_{nm}}p_{\xi}(j|x_{2})\cdot\mathds{1}\{x_{1}=x_{1}^{(nm)}[j]\}.$ (12)
@@
$\xi^{*}=\arg\max_{\xi}\sum_{t\in D_{x}}w_{t}\log p_{\xi}(j_{t}|x_{2,t}).$ (16)
When $x_{1,i}$ is missing, the sampling is enriched by simulated $x_{1,i,r}$. Eq. 16 selects $j^{*}$ from $I_{nm}$ based on $p_{\xi^{\prime}}(j|x_{2})$ according to Eq. 12 and Eq. 11. The $D_{x}$ above is a subset of the data for $\xi$, learned on the missing data part. We also note that learning the parameters above is parallelizable.

### v4

@@ Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals with any ML model. SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation. Through extensive empirical evaluation of diverse simulated distributions and 11 real-world tabular datasets, SEMF consistently produces narrower prediction intervals while maintaining the desired coverage probability, outperforming traditional quantile regression methods. Furthermore, without using the quantile (pinball) loss, SEMF allows point predictors, including gradient-boosted trees and neural networks, to be calibrated with conformal quantile regression. The results indicate that SEMF enhances uncertainty quantification under diverse data distributions and is particularly effective for models that otherwise struggle with inherent uncertainty representation. [^1] Code: https://github.com/Unco3892/semf-paper
@@ 1 Introduction
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic), with a particular focus on uncertainty quantification. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. The EM algorithm has already been used for supervised learning tasks using specific models [8, 9, 10], minimal supervision [11], and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 13]. Our method, SEMF, differs from prominent supervised EM approaches such as ghahramani1993em [8], which focus on point prediction using Gaussian Mixture Models (GMMs). Furthermore, SEMF operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions. Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper. Finally, SEMF generates representations for latent modalities through specialized models, holding potential for multi-modal data applications, though multi-modal settings are left for future exploration.
The remainder of this paper is organized as follows: Section 2 details the theory and methodology of SEMF. Section 3 reviews related work in latent representation learning and uncertainty estimation. Section 4 describes the experimental setup, including synthetic and benchmark datasets, as well as evaluation metrics. Section 5 discusses the results, demonstrating the efficacy of SEMF. Lastly, Section 6 concludes the paper, and Section 7 outlines the limitations and potential research directions.
@@ 2 Method
This section presents the founding principles of SEMF from its parameters, training, and inference procedure with, at its core, the EM algorithm. Invented to maximize the model likelihood, it builds a sequence of parameters that guarantee an increase in the log-likelihood [14] by iterating between the Expectation (E) and the Maximization (M) steps. In the E-step, one computes
@@
where $p^{\prime}$ stands for the current estimates, $\log p(x,z)$ is the log-likelihood of the complete observation $(x,z)$, and $z$ is a latent variable. The M-step updates the current estimates with the arguments maximizing the $Q$-function: $p^{\prime}\leftarrow\arg\max_{p}Q(p|p^{\prime})$ The sequence is repeated until convergence.
@@ 2.1 Problem Scenario
Let $x=(x_{1},x_{2},\ldots,x_{K})$ denote $K$ inputs and the output be $y$. For simplicity, we limit $y$ to be continuous numerical, although this assumption could be relaxed to discrete or categorical without loss of generality. Component $x_{k}$ is a source: a modality, a single or group of variables, or an unstructured input such as an image or text. For clarity, we limit to $K=2$, where $x_{1}$ and $x_{2}$ are single variables.
@@
The contribution to the log-likelihood of a complete observation $(y,z,x)$ is $\log p(y,z|x)=\log p(y|z)+\log p(z|x)$. In the E-step, we compute
@@
Eq. 2 can be estimated by MC sampling. Since sampling from $p^{\prime}(z|y,x)$ can be inefficient, we rather rely on the decomposition $p^{\prime}(z|y,x)=p^{\prime}(y|z)p^{\prime}(z|x)/p^{\prime}(y|x)$. Thus, we sample $z_{r}$ from $p^{\prime}(z|x)$, $r=1,\ldots,R$, and, setting $w_{r}=p^{\prime}(y|z_{r})/\sum_{t}p^{\prime}(y|z_{t})$, approximate the right-hand side term of Eq. 2
@@ 2.2 Objective Function
Adapting Eq. 3 for the observed data $\{(y_{i},x_{i})\}_{i=1}^{N}$, the overall loss function, $\mathcal{L}$, is
$\mathcal{L}(\phi,\theta)=-\sum_{i=1}^{N}\sum_{r=1}^{R}\left\{\log p_{\phi}(z_{i,r}|x_{i})+\log p_{\theta}(y_{i}|z_{i,r})\right\}w_{i,r},$ (4)
where the models of $p(y|z)$ and $p(z|x)$ inherit parameters $\theta$ and $\phi$, respectively. The weights are
$w_{i,r}=\frac{p_{\theta^{\prime}}(y_{i}|z_{i,r})}{\sum_{t=1}^{R}p_{\theta^{\prime}}(y_{i}|z_{i,t})}$ (5)
where $z_{i},r\stackrel{{\scriptstyle ind.}}{{\sim}}p_{\phi^{\prime}}(z|x_{i})$, $r=1,\ldots,R$.
Eq. 4 shows that ${\cal L}$ is a sum of losses associated with the encoder model, $p_{\phi}$, for each source, and the decoder model, $p_{\theta}$, from the latent variables to the output. $p_{\theta}$ and $p_{\phi}$ are referred to as encoder and decoder in the spirit of auto-encoder models. At each M-step, ${\cal L}$ is minimized with respect to $\theta$ and $\phi$. Then, $\theta^{\prime}$ and $\phi^{\prime}$ are updated, as well as the weights and the sampled $z$. Then, the process is iterated until convergence.
2.2.1 Example: $\mathcal{L}(\phi,\theta)$ under normality
For illustration purposes, we develop below the case where $p_{\phi}$ and $p_{\theta}$ are normal distributions, similar to kingma2014vae [7]. As a reminder, any other distributions could be adopted, including non-continuous or non-numerical outcomes.
@@ 2.2.1 Example: $\mathcal{L}(\phi,\theta)$ under normality
$\log p_{\phi}(z_{k}|x_{k})$ $=$ $-\frac{m_{k}}{2}\log 2\pi-\frac{m_{k}}{2}\log\sigma_{k}^{2}-\frac{1}{2\sigma_{k}^{2}}\sum_{j=1}^{m_{k}}\{z_{k,j}-g_{\phi_{k},j}(x_{k})\}^{2},\quad k=1,2.$ (7)
The mean $g_{\phi_{k}}(x_{k})$ can be any model, such as a neural network, with output of length $m_{k}$, $k=1,2$. The scale $\sigma_{k}$ can be fixed, computed via the weighted residuals, or learned through a separate set of models. It controls the amount of noise introduced in the latent dimension and is pivotal in determining the prediction interval width for $p(y|z)$. The code implementation allows for all three methods. This paper only presents the results for a fixed $\sigma_{k}$, as well as training a separate set of $g_{\phi_{k}}(x_{k})$ models to estimate varying $\sigma_{k}$s.